# Simulation studies of the isovector reorientation effect of deuteron scattering on heavy target

### v2

@@ Abstract
The isovector reorientation (IVR) effect of deuteron scattering on heavy target provides a novel means to probe the nuclear isovector potential, which gives rise to the nuclear symmetry energy. The simulation studies on the experimental measurement of IVR effect using the SAMURAI terminal at RIKEN Nishina center have been performed to demonstrate the feasibility of the experiment. By introducing a well-designed polarimeter to detect the $\mathrm{p}(\vec{\mathrm{d}},\mathrm{d})\mathrm{p}$ elastic scattering, monitoring of the tensor polarization of the deuteron beam can be implemented. The protons and neutrons produced by the breakup of polarized deuterons scattering off heavy targets are designed to be measured by proton drift chamber (PDC) combined with the SAMURAI magnet and NEBULA detector, respectively. The detector responses are simulated using Geant4 framework, where the events of the deuteron elastic breakup are generated by an Improved Quantum Molecular Dynamics model. The results of reconstructing the deuteron breakup events demonstrate the feasibility of detecting the IVR effect at SAMURAI with both longitudinal and transverse tensor polarized deuteron beams with a polarization degree of approximately 80%.
Analysing Power; Isovector Reorientation Effect; Tensor polarization; ImQMD
@@ I Introduction
The symmetry energy characterizes the isospin-dependent part of the nuclear equation of state, quantifying the energy cost associated with neutron-proton asymmetry Li et al. (2008); Danielewicz et al. (2002); Li et al. (2021). The density dependence of the symmetry energy, $E_{\rm sym}(\rho)$, has drawn increasing attention from both nuclear physics and astrophysics communities. However, the stiffness of $E_{\rm sym}(\rho)$  still suffers from significant uncertainty, particularly at supra-saturation densities Li et al. (2019). Recently, enormous progress has been made from heavy ion experiments Tsang et al. (2004); Estee et al. (2021); Xiao et al. (2009); Zhang et al. (2017) and observations of gravitational waves generated by neutron star merger Abbott et al. (2017, 2018), as well as the combination of the two Tsang et al. (2024); Huth et al. (2022). Further elaborations are still ongoing in searching new probes of $E_{\rm sym}(\rho)$Wang et al. (2022, 2023a, 2023b).
@@
$=\frac{k_{\rm F}^{2}}{6m^{*}_{0}(\rho,k_{\rm F})}+\frac{1}{2}U_{\rm sym}(\rho,k_{\rm F})$
where $t(k_{\mathrm{F}})$ is the nucleon kinetic energy at the Fermi momentum $k_{\mathrm{F}}=\left(\frac{3\pi^{2}\rho}{2}\right)^{1/3}$ at the symmetric nuclear matter density $\rho$. Equation I shows that the microscopic origin of the symmetry energy is the difference between the neutron and proton Fermi surfaces. For the isoscalar potential $U_{0}(\rho,k)$, information on its density and momentum dependence has been obtained from high-energy heavy-ion collisions (e.g., see Ref. Danielewicz et al. (2002)). While existing data constrain its behavior to some extent, significant uncertainties persist in high-momentum and high-density regimes. In contrast, the isovector potential $U_{\mathrm{sym}}(\rho,k)$, particularly at extreme densities and momenta, remains poorly constrained. This potential has been identified as a critical parameter for resolving ambiguities in the high-density behavior of the symmetry energy Xu et al. (2011); Chen et al. (2012); Li et al. (2008). At saturation density, the isoscalar effective mass $\frac{m_{0}^{*}}{m}=0.70\pm 0.05$, and the isovector potential $U_{\rm sym}(\rho_{0},k_{\rm F})=28.7\pm 7.8~{}\mathrm{MeV}$ Li and Han (2013); Wang et al. (2023c). Therefore, the challenge of excluding the interference of isoscalar forces in heavy-ion experiments while extracting the characteristics of isovector forces has become critical.
The isovector reorientation (IVR) effect effectively mitigates the influence of the isoscalar potential, enabling the extraction of isovector information with high sensitivity Ou et al. (2 = 15); Liang et al. (2020). As shown in Fig. 1, in the peripheral scattering of a polarized deuteron beam on a heavy target, the isovector force, which is attractive to protons and repulsive to neutrons, acts like a torque on the deuteron, inducing an additional rotation of the latter. Consequently, the subsequent breakup incorporates the rotational effect, which exhibits high sensitivity to the isovector force. Detailed simulations demonstrate that the angular distribution of the correlated neutron and proton from the deuteron breakup is highly sensitive to the isovector potential, yet insensitive to the isoscalar nuclear potential. The experimental observable is straightforward to implement: measuring the angular distribution of the relative momentum vector between neutrons and protons arising from the breakup of a polarized deuteron. This observable acts as a probe with exceptional sensitivity to the symmetry energy in regions of low nuclear density. By employing this probe, the uncertainties in the density dependence of the symmetry energy can be effectively constrained.
@@
Among various research and development (R&D) works, two key issues have to be addressed before conducting the experiment at SAMURAI, RIKEN Nishina center. One is the monitoring of the polarized deuteron beam, the other is the design (and optimization) of the detector setup to measure the breakup of the polarized deuteron scattering on heavy target, which carries the IVR effect. These two issues can be elucidated by simulation studies, which make the current paper. The remainder of this paper is organized as following. In Section II, we outline the experimental plan and describe the simulation methods employed in this study. Section III focuses on the monitoring of the polarized beam, where we define tensor polarization and introduce two monitoring options. Moreover, the design of the polarimeter and the simulation results for beam monitoring are presented. Section IV describes the ImQMD transport model as the event generator in our simulations. Section V present the simulation of the detector response at SAMURAI and the event reconstruction. We detail the detector setup, target position optimization, and the responses of both neutrons and protons, along with the background simulation and the beam dumper. Additionally, we observe the IVR effect within the current experimental design. Finally, Section VI provides the conclusion and outlooks of the study.
@@ II A brief introduction to the beam experiment
The experimental setup is specifically optimized to measure the isovector reorientation effect in deuteron scattering on heavy targets. Figure 2 presents the overview of the beam line and detector arrangement, highlighting the strategic placement of key components to maximize detection efficiency and minimize background. The deuteron beam is polarized and accelerated to 190 MeV/u. The tensor polarized beam is provided by the polarized ion source and accelerated by the AVF and RRC successively, and then by SRC through the IRC bypass transport beam line. The spin symmetry axis of the deuteron beams is controlled by the Wien Filter prior to acceleration. Single-turn extractions are required for the three cyclotrons to maintain polarization amplitudes during acceleration. The polarization of the deuteron beam is typically 80% of the theoretical valueSakamoto et al. (2016).
Figure 2: The bird’s-eye view of the experiment.
@@
We intend to optimize the reaction vertex in the SAMURAI dipole to make the best use of the coverage of the detectors. According to the Improved Quantum Molecular Dynamics (ImQMD) model calculations, for the scattering events with $b=7\sim 9\,\mathrm{fm}$ which is the range showing the most sensitivity of IVR effect on the $\gamma$ parameter, which is introduced to mimic different density dependence of $E_{\rm sym}(\rho)$, the most probable direction of the breakup neutrons is not very close to $\theta=0{{}^{\circ}}$. Considering that NEBULA is difficult to move, in order to measure the neutrons in the most favorable range of $\theta_{\mathrm{n}}$, it is suggested to place the target upstream of the magnet, at a distance of 3 meters from the center of the magnet, which ensures that the range of $-10{{}^{\circ}}\leq\theta_{\mathrm{n}}\leq 10{{}^{\circ}}$ is covered for the most probable neutron emissions.
As for the SAMURAI magnet, we choose $B=1.2\,\mathrm{T}$ and the $30{{}^{\circ}}$ configuration, which fits mostly our experimental purpose. In addition, we do not require the magnet chamber to be in vacuum, since the beam trajectory is influenced little by the air at the current beam energy. By removing the downstream exit window of the SAMURAI chamber, we have a wider range of exit angle for both neutrons and protons. According to the simulation results, the reorientation effect can be observed for both $z$-polarization ($p_{zz}$) and $y$-polarization ($p_{yy}$). Therefore, we prefer to use either the $z$-polarization or $y$-polarization in our experiment, depending on the machine condition. The unreacted deuteron beam will be collected by the beam dumper, which is surrounded by water tank to shield the neutron background radiation.
To achieve the experimental goal, the detector system is arranged with a polarimeter and a veto detector upstream of the target to monitor deuteron polarization and suppress inelastic background, respectively. Downstream of the target, protons from deuteron breakup are detected by the proton drift chamber (PDC) backed by a time-of-flight (TOF) array of plastic scintillators, while neutrons are detected by the NEBULA array. The PDC is positioned 3 meters from the target, and NEBULA is placed 7.3 meters from the target, ensuring optimal detection efficiency for both protons and neutrons.
@@ III Monitor the tensor polarization
For this experiment, a polarized deuteron beam is essential. Because of the unpolarized deuteron beam, the random orientation will average out the differences in the proton-neutron angular distribution.
To ensure accurate measurements, we monitor the deuteron beam’s polarization using a method based on d-p elastic scattering. This involves measuring the cross-sections at different angles to confirm the polarization state. While the theoretical framework for polarized nuclear reactions is well established and experimental techniques for monitoring $p_{y^{\prime}y^{\prime}}$ are available, the direct measurement of $p_{z^{\prime}z^{\prime}}$ remains unexplored. This work addresses this gap by developing a robust monitoring mechanism for $p_{z^{\prime}z^{\prime}}$. To avoid misunderstandings caused by different notational conventions, we start from the most fundamental principles and derive our polarization monitoring mechanism in detail.
III.1 Defination of the tensor polarization
Deuteron is a spin-1 nucleus. The polarization state of the deuteron can be fully described using its spin density matrix. This matrix characterizes the statistical distribution of the mixed quantum states $\rho=\sum_{i}{Q}_{i}\ket*{\phi}\bra*{\phi}$ |. For a spin-1 system such as the deuteron, the density matrix $\rho$ is a $3\times 3$ Hermitian matrix that encodes the information about both the vector and the tensor polarization, the latter of which is required for this experiment.
Here we employ the Cartesian basis for the expansion, consistent with the convention in reference Ohlsen (1972), as it aligns with the definition of the tensor polarization of the deuteron beam used in the experiment along the z and y directions. Although Cartesian tensors suffer from the reducibility difficulty, i.e., they can be decomposed into three independent components with different transformation properties under rotation corresponding to angular momentum multiplicities $L=0$, 1, and 2, respectively. Cartesian tensors can be decomposed into three tensors that transform according to the spherical harmonics of order 0, 1, and 2. Therefore, spherical tensors are more fundamental and better keep the symmetry. Cartesian and spherical bases can be readily transformed into each other, as detailed in references Ohlsen (1972) and Rose (1995). Specifically, for the deuteron with spin $S=1$, the vector polarization is given by $\mathscr{P}_{i}=S_{i}$, and the tensor operator is expressed as $\mathscr{P}_{ij}=3S_{i}S_{j}-2I$, where $S_{i}$ and $S_{j}$ are the spin operators and $I$ is the identity matrix. Consequently, the density matrix can be expanded in terms of these operators.
$\rho=\frac{1}{3}\Big{\{}I+\frac{3}{2}\left(p_{x}\mathscr{P}_{x}+p_{y}\mathscr{P}_{y}+p_{z}\mathscr{P}_{z}\right)$
@@ III.1 Defination of the tensor polarization
$+\frac{1}{3}\left(p_{xx}\mathscr{P}_{xx}+p_{yy}\mathscr{P}_{yy}+p_{zz}\mathscr{P}_{zz}\right)\Big{\}}.$ (2)
Where the terms $p_{i}$ and $p_{ij}$ represent the vector and tensor polarization components of the initial deuteron beam, respectively. However, $\mathscr{P}_{xx}$,$\mathscr{P}_{yy}$,$\mathscr{P}_{zz}$ is linear dependent. Based on $\mathscr{P}_{xx}+\mathscr{P}_{yy}+\mathscr{P}_{zz}=0$, we can rewrite the density matrix as other forms.
@@ III.2 Monitor method
$=\sigma_{0}(\theta)\left\{1+\frac{3}{2}\sum_{i}p_{i}A_{i}+\frac{1}{3}\sum_{ij}p_{ij}A_{ij}\right\}$ (7)
where $\sigma_{0}(\theta)=\frac{1}{3}{\rm Tr}M\rho_{\rm unpol}M^{\dagger}=\frac{1}{3}{\rm Tr}MM^{\dagger}$. Here, $\sigma$ is the differential cross-section for a polarized beam, $M$ is the scattering matrix describing the reaction, and $M^{\dagger}$ is its Hermitian conjugate. The terms $p_{i}$ and $p_{ij}$ represent the vector and tensor polarization components of the initial deuteron beam. Here, $p$ corresponds to the statistical probability of the polarization operator $\mathscr{P}$, which is equivalent to the $Q$ defined earlier. $A_{i}$ and $A_{ij}$ are the corresponding vector and tensor analyzing powers, which characterize the sensitivity of the reaction to the polarization. $\sigma_{0}(\theta)$ is the unpolarized differential cross-section at scattering angle $\theta$. This formula shows how the measured cross-section depends on both the polarization state of the beam and the analyzing powers of the reaction.
$x^{\prime}$$y^{\prime}$$z^{\prime}$$x$$y$$z$$beam$$\phi$
Figure 3: Reference frame. The red axes represent the projectile helicity frame, with the $xoz$ plane defined as the reaction plane.
@@
$\left[A_{xx}(\theta)-A_{yy}(\theta)\right]+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}(\theta)$ (8)
@@
$\sigma_{L}$ $=\sigma_{0}\left\{1+\frac{3}{2}p_{y^{\prime}}A_{y}+\frac{2}{3}p_{x^{\prime}z^{\prime}}A_{xz}+\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\}$ (9)
$\sigma_{R}$ $=\sigma_{0}\left\{1-\frac{3}{2}p_{y^{\prime}}A_{y}-\frac{2}{3}p_{x^{\prime}z^{\prime}}A_{xz}+\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\}$
$\sigma_{U}$ $=\sigma_{0}\left\{1-\frac{3}{2}p_{x^{\prime}}A_{y}+\frac{2}{3}p_{y^{\prime}z^{\prime}}A_{xz}-\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\}$
$\sigma_{D}$ $=\sigma_{0}\left\{1+\frac{3}{2}p_{x^{\prime}}A_{y}-\frac{2}{3}p_{y^{\prime}z^{\prime}}A_{xz}-\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\}$
In reference Bieber et al. (2001), the method for monitoring $P_{y^{\prime}y^{\prime}}$ has been described. (In the paper Bieber et al. (2001), it should be noted that the definition of the z-direction for the deuteron polarization differs from the one used in the schematic diagram of Fig. 7 therein. Additionally, their equation (5) only considers the polarization axis is along $y^{\prime}$, and $T_{22}\equiv\operatorname{Re}T_{22}$ due to the parity conservation.) With the present notation used here, the equation (5) in Bieber et al. (2001) can be written as
[…]
For the monitoring of $p_{z^{\prime}z^{\prime}}$, we define the average differential cross section of the four detectors as $\bar{\sigma}=\frac{\sigma_{L}+\sigma_{R}+\sigma_{D}+\sigma_{U}}{4}$. The relationship between the average intensity and the tensor polarization is given by Eq. (11). If we know the analyzing power, we can determine the tensor polarization from the average intensity. Alternatively, we can take measurements at two different angles to eliminate the effects of beam intensity and other factors.
$\bar{\sigma}\equiv\frac{\sigma_{L}+\sigma_{R}+\sigma_{D}+\sigma_{U}}{4}=\sigma_{0}\left(1+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right)$ (11)

### v3

@@ Abstract
The isovector reorientation (IVR) effect of deuteron scattering on heavy target provides a novel means to probe the nuclear isovector potential, which gives rise to the nuclear symmetry energy. The simulation studies on the experimental measurement of IVR effect using the SAMURAI terminal at RIKEN Nishina center have been performed to demonstrate the feasibility of the experiment. By introducing a well-designed polarimeter to detect the $\mathrm{p}(\vec{\mathrm{d}},\mathrm{d})\mathrm{p}$ elastic scattering, monitoring of the tensor polarization of the deuteron beam can be implemented. The protons and neutrons produced by the breakup of polarized deuterons scattering off heavy targets are designed to be measured by proton drift chamber (PDC) combined with the SAMURAI magnet and NEBULA (Neutron-detection system for Breakup of Unstable-Nuclei with Large Acceptance) detector, respectively. The detector responses are simulated using GEANT4 framework, where the events of the deuteron elastic breakup are generated by an Improved Quantum Molecular Dynamics model. The results of reconstructing the deuteron breakup events demonstrate the feasibility of detecting the IVR effect at SAMURAI with both longitudinal and transverse tensor polarized deuteron beams with a polarization degree of approximately 80%.
Analysing Power; Isovector Reorientation Effect; Tensor polarization; ImQMD; Deuteron-induced Reaction
@@ I Introduction
The symmetry energy characterizes the isospin-dependent part of the nuclear equation of state, quantifying the energy cost associated with neutron-proton asymmetry Li et al. (2008); Danielewicz et al. (2002); Li et al. (2021). The density dependence of the symmetry energy, $E_{\rm sym}(\rho)$, has drawn increasing attention from both nuclear physics and astrophysics communities. However, the stiffness of $E_{\rm sym}(\rho)$  still suffers from significant uncertainty, particularly at supra-saturation densities Li et al. (2019). Recently, enormous progress has been made from heavy ion experiments Tsang et al. (2004); Estee et al. (2021); Xiao et al. (2009); Zhang et al. (2017) and observations of gravitational waves generated by neutron star merger Abbott et al. (2017, 2018), as well as the combination of the two Tsang et al. (2024); Huth et al. (2022). Further elaborations are still ongoing in searching new probes of $E_{\rm sym}(\rho)$Wang et al. (2022, 2023a, 2025).
@@
$=\frac{k_{\rm F}^{2}}{6m^{*}_{0}(\rho,k_{\rm F})}+\frac{1}{2}U_{\rm sym}(\rho,k_{\rm F}),$
Here, $t(k_{\mathrm{F}})$ is the nucleon kinetic energy at the Fermi momentum, and $k_{\mathrm{F}}=\left(\frac{3\pi^{2}\rho}{2}\right)^{1/3}$ is the Fermi momentum for symmetric nuclear matter with density $\rho$. Equation I shows that the microscopic origin of the symmetry energy is the difference between the neutron and proton Fermi surfaces. Equation I shows that the microscopic origin of the symmetry energy is the difference between the neutron and proton Fermi surfaces. For the isoscalar potential $U_{0}(\rho,k)$, information on its density and momentum dependence has been obtained from high-energy heavy-ion collisions (e.g., see Ref. Danielewicz et al. (2002)). While existing data constrain its behavior to some extent, significant uncertainties persist in high-momentum and high-density regimes. In contrast, the isovector potential $U_{\mathrm{sym}}(\rho,k)$, particularly at extreme densities and momenta, remains poorly constrained. This potential has been identified as a critical parameter for resolving ambiguities in the high-density behavior of the symmetry energy Xu et al. (2011); Chen et al. (2012); Li et al. (2008). At saturation density, the isoscalar effective mass $\frac{m_{0}^{*}}{m}=0.70\pm 0.05$, and the isovector potential $U_{\rm sym}(\rho_{0},k_{\rm F})=28.7\pm 7.8~\mathrm{MeV}$ Li and Han (2013); Wang et al. (2023b). Therefore, the challenge of excluding the interference of isoscalar forces in heavy-ion experiments while extracting the characteristics of isovector forces has become critical.
The isovector reorientation (IVR) effect effectively mitigates the influence of the isoscalar potential, enabling the extraction of isovector information with high sensitivity Ou et al. (2 = 15); Liang et al. (2020). As shown in Figure 1, in the peripheral scattering of a polarized deuteron beam on a heavy target, the isovector force, which is attractive to protons and repulsive to neutrons, acts like a torque on the deuteron, inducing an additional rotation of the latter. Consequently, the subsequent breakup incorporates the rotational effect, which exhibits high sensitivity to the isovector force. Detailed simulations demonstrate that the angular distribution of the correlated neutron and proton from the deuteron breakup is highly sensitive to the isovector potential, yet insensitive to the isoscalar nuclear potential. The experimental observable is straightforward to implement: measuring the angular distribution of the relative momentum vector between neutrons and protons arising from the breakup of a polarized deuteron. This observable acts as a probe with exceptional sensitivity to the symmetry energy in regions of low nuclear density. By employing this probe, the uncertainties in the density dependence of the symmetry energy can be effectively constrained.
@@
Among various research and development (R&D) works, two key issues have to be addressed before conducting the experiment at SAMURAI, RIKEN Nishina center. One is the monitoring of the polarized deuteron beam, the other is the design (and optimization) of the detector setup to measure the breakup of the polarized deuteron scattering on heavy target, which carries the IVR effect. These two issues can be elucidated by simulation studies, which make the current paper. The remainder of this paper is organized as following. In Section II, we outline the experimental plan and describe the simulation methods employed in this study. Section III focuses on the monitoring of the polarized beam, where we define tensor polarization and introduce two monitoring options. Moreover, the design of the polarimeter and the simulation results for beam monitoring are presented. Section IV describes the Improved Quantum Molecular Dynamics(ImQMD) transport model as the event generator in our simulations. Section V present the simulation of the detector response at SAMURAI and the event reconstruction. We detail the detector setup, target position optimization, and the responses of both neutrons and protons, along with the background simulation and the beam dumper. Additionally, we observe the IVR effect within the current experimental design. Finally, Section VI provides the conclusion and outlooks of the study.
@@ II A brief introduction to the beam experiment
The experimental setup is specifically optimized to measure the isovector reorientation effect in deuteron scattering on heavy targets. Figure 2 presents the overview of the beam line and detector arrangement, highlighting the strategic placement of key components to maximize detection efficiency and minimize background. The deuteron beam is polarized and accelerated to 190 MeV/u. The tensor polarized beam is provided by the polarized ion source and accelerated by the AVF (Azimuthally Varying Field Ring Cyclotron) and RRC (Riken Ring Cyclotron) successively, and then by SRC (Superconducting Ring Cyclotron) through the IRC (Intermediate-stage Ring Cyclotron) bypass transport beam line. The spin symmetry axis of the deuteron beams is controlled by the Wien Filter prior to acceleration. Single-turn extractions are required for the three cyclotrons to maintain polarization amplitudes during acceleration. The polarization of the deuteron beam is typically 80% of the theoretical value Sakamoto et al. (2016).
Figure 2: The bird’s-eye view of the experiment. The AVF (Azimuthally Varying Field Ring Cyclotron), RRC (Riken Ring Cyclotron) and SRC (Superconducting Ring Cyclotron) are used to produce, accelerate and transport the polarized deuteron beam to 190 MeV/u. Beam trigger (SBT) is to provide the ount number of beam and determine start timing to calculate the time of flight(TOF). The Polarimeter monitors the tensor polarization of the deuteron beam before it reaches the target. The target is positioned at the entrance of the SAMURAI spectrometer, which is a large-acceptance superconducting dipole magnet for momentum analysis of charged particles. The PDC (Proton Drift Chamber) detects breakup protons downstream of the target, while NEBULA is used to detect neutrons.The VD (veto detector) is to veto the central collision based on the multiplicity.
@@
The experimental setup is specifically optimized to mea- sure the isovector reorientation effect in deuteron scattering on heavy targets. According to the Improved Quantum Molecular Dynamics (ImQMD) model calculations, for the scattering events with impact parameter $b=7\sim 9\,\mathrm{fm}$ which is the range showing the most sensitivity of IVR effect on the stiffness of the symmetry energy $E_{\rm sym}(\rho)$, as characterized by the parameter of $\gamma$ in Eq. (14), the most probable direction of the breakup neutrons is not very close to $\theta=0{{}^{\circ}}$. Considering that NEBULA is difficult to move, in order to measure the neutrons in the most favorable range of $\theta_{\mathrm{n}}$, it is suggested to place the target upstream of the magnet, at a distance of 3 meters from the center of the magnet, which ensures that the range of $10{{}^{\circ}}\leq\theta_{\mathrm{n}}\leq 10{{}^{\circ}}$ is covered for the most probable neutron emissions.
As for the SAMURAI magnet, we choose $B=1.2\,\mathrm{T}$ and the $30{{}^{\circ}}$ configuration, which fits mostly our experimental purpose. In addition, we do not require the magnet chamber to be in vacuum, since the beam trajectory is influenced little by the air at the current beam energy. By removing the downstream exit window of the SAMURAI chamber, we have a wider range of exit angle for both neutrons and protons. According to the simulation results, the reorientation effect can be observed for both $z$-polarization ($p_{zz}$) and $y$-polarization ($p_{yy}$). Therefore, we prefer to use either the $z$-polarization or $y$-polarization in our experiment, depending on the machine condition. The unreacted deuteron beam will be collected by the beam dumper, which is surrounded by a water tank to shield the neutron background radiation.
To achieve the experimental goal, the detector system is arranged with a polarimeter and a veto detector upstream of the target to monitor deuteron polarization and suppress inelastic background, respectively. Downstream of the target, protons from deuteron breakup are detected by the proton drift chamber (PDC), while neutrons are detected by the NEBULA array. The PDC is positioned 3 meters from the target, and NEBULA is placed 7.3 meters from the target, ensuring optimal detection efficiency for both protons and neutrons.
@@ III Monitor the tensor polarization
The design with a polarized deuteron beam is essential in this experiment. If the deuteron beam were unpolarized, the random orientation would average out the differences in the proton-neutron angular distribution, making it impossible to observe the IVR effect. We monitor the deuteron beam’s polarization using a method based on d-p elastic scattering. This involves measuring the cross-sections at different angles to confirm the polarization state. While the theoretical framework for polarized nuclear reactions is well established and experimental techniques for monitoring $p_{y^{\prime}y^{\prime}}$ are available, the direct measurement of $p_{z^{\prime}z^{\prime}}$ remains unexplored. This work addresses this gap by developing a robust monitoring mechanism for $p_{z^{\prime}z^{\prime}}$. To avoid misunderstandings caused by different notational conventions, we start from the most fundamental principles and derive our polarization monitoring mechanism in detail.
III.1 Definition of the tensor polarization
Deuteron is a spin-1 nucleus. The polarization state of the deuteron can be fully described using its spin density operator. This operator characterizes a statistical mixture of pure quantum states,
$|,\displaystyle\hat{\rho}=\sum_{i}Q_{i}\ket{\phi_{i}}\bra{\phi_{i}}$, (2)
[…]
Here we employ the Cartesian basis for the expansion, consistent with the convention in reference Ohlsen (1972), as it aligns with the definition of the tensor polarization of the deuteron beam used in the experiment along the z and y directions. Although Cartesian tensors suffer from the reducibility difficulty, i.e., they can be decomposed into three independent components with different transformation properties under rotation corresponding to angular momentum multiplicities $L=0$, 1, and 2, respectively. Cartesian tensors can be decomposed into three tensors that transform according to the spherical harmonics of order 0, 1, and 2. Therefore, spherical tensors are more fundamental and better preserve the symmetry. Cartesian and spherical bases can be readily transformed into each other, as detailed in references Ohlsen (1972) and Rose (1995). Specifically, for the deuteron with spin $S=1$, the vector polarization is given by $\mathscr{P}_{i}=S_{i}$, and the tensor operator is expressed as $\mathscr{P}_{ij}=3S_{i}S_{j}-2I$, where $S_{i}$ and $S_{j}$ are the spin operators and $I$ is the identity matrix. Consequently, the density matrix can be expanded in terms of these operators.
$\hat{\rho}=\frac{1}{3}\Big\{I+\frac{3}{2}\left(p_{x}\mathscr{P}_{x}+p_{y}\mathscr{P}_{y}+p_{z}\mathscr{P}_{z}\right)$
@@ III.1 Definition of the tensor polarization
$+\frac{1}{3}\left(p_{xx}\mathscr{P}_{xx}+p_{yy}\mathscr{P}_{yy}+p_{zz}\mathscr{P}_{zz}\right)\Big\},$ (3)
Where the terms $p_{i}$ and $p_{ij}$ represent the vector and tensor polarization components of the initial deuteron beam, respectively.
However, $\mathscr{P}_{xx}$$\mathscr{P}_{yy}$$\mathscr{P}_{zz}$ is linear dependent. Based on $\mathscr{P}_{xx}+\mathscr{P}_{yy}+\mathscr{P}_{zz}=0$, we can rewrite the density matrix as other forms.
@@ III.2 Monitor method
$=\sigma_{0}(\theta)\left\{1+\frac{3}{2}\sum_{i}p_{i}A_{i}+\frac{1}{3}\sum_{ij}p_{ij}A_{ij}\right\},$ (8)
where $\sigma_{0}(\theta)=\frac{1}{3}{\rm Tr}M\hat{\rho}_{\rm unpol}M^{\dagger}=\frac{1}{3}{\rm Tr}MM^{\dagger}$. Here, $\sigma$ is the differential cross-section for a polarized beam, $M$ is the scattering matrix describing the reaction, and $M^{\dagger}$ is its Hermitian conjugate. The terms $p_{i}$ and $p_{ij}$ represent the vector and tensor polarization components of the initial deuteron beam. Here, $p$ corresponds to the statistical probability of the polarization operator $\mathscr{P}$, which is equivalent to the $Q$ defined earlier. $A_{i}$ and $A_{ij}$ are the corresponding vector and tensor analyzing powers, which characterize the sensitivity of the reaction to the polarization. $\sigma_{0}(\theta)$ is the unpolarized differential cross-section at scattering angle $\theta$. This formula shows how the measured cross-section depends on both the polarization state of the beam and the analyzing powers of the reaction.
[…]
Figure 3: Reference frame. The red axes represent the projectile helicity frame, with the xozxoz plane defined as the reaction plane.
@@
$\left[A_{xx}(\theta)-A_{yy}(\theta)\right]+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}(\theta).$ (9)
@@
$\sigma_{L}$ $=\sigma_{0}\left\{1+\frac{3}{2}p_{y^{\prime}}A_{y}+\frac{2}{3}p_{x^{\prime}z^{\prime}}A_{xz}+\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\},$ (10)
$\sigma_{R}$ $=\sigma_{0}\left\{1-\frac{3}{2}p_{y^{\prime}}A_{y}-\frac{2}{3}p_{x^{\prime}z^{\prime}}A_{xz}+\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\},$
$\sigma_{U}$ $=\sigma_{0}\left\{1-\frac{3}{2}p_{x^{\prime}}A_{y}+\frac{2}{3}p_{y^{\prime}z^{\prime}}A_{xz}-\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\},$
$\sigma_{D}$ $=\sigma_{0}\left\{1+\frac{3}{2}p_{x^{\prime}}A_{y}-\frac{2}{3}p_{y^{\prime}z^{\prime}}A_{xz}-\frac{1}{6}\left(p_{x^{\prime}x^{\prime}}-p_{y^{\prime}y^{\prime}}\right)\left(A_{xx}-A_{yy}\right)+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right\}.$
In paper Bieber et al. (2001) the method to monitor $p_{y^{\prime}y^{\prime}}$ is described. Note that their axis convention for the deuteron polarization differs from the convention used here, and their Eq (5) assumes the polarization axis along $y^{\prime}$. Using the notation adopted in this work, one first defines the measurable count asymmetry
$\begin{aligned} R_{LRUD}&=\frac{N_{L}+N_{R}-N_{U}-N_{D}}{N_{L}+N_{R}+N_{U}+N_{D}}=\frac{p_{y^{\prime}y^{\prime}}(A_{xx}-A_{yy})}{2p_{y^{\prime}y^{\prime}}A_{zz}-4},\\ p_{y^{\prime}y^{\prime}}&=\frac{R_{LRUD}}{\frac{1}{2}A_{zz}R_{LRUD}-\frac{1}{4}(A_{xx}-A_{yy})}.\end{aligned}$ (11)
For the monitoring of $p_{z^{\prime}z^{\prime}}$, we define the average differential cross section of the four detectors as $\bar{\sigma}=\frac{\sigma_{L}+\sigma_{R}+\sigma_{D}+\sigma_{U}}{4}$. The relationship between the average intensity and the tensor polarization is given by Eq. (12). If we know the analyzing power, we can determine the tensor polarization from the average intensity. Alternatively, we can take measurements at two different angles to eliminate the effects of beam intensity and other factors:
$\bar{\sigma}\equiv\frac{\sigma_{L}+\sigma_{R}+\sigma_{D}+\sigma_{U}}{4}=\sigma_{0}\left(1+\frac{1}{2}p_{z^{\prime}z^{\prime}}A_{zz}\right).$ (12)